OpenStudy (adi3):

Pls help The cyclist is applying a forward push with pedals of 80N, while there is a backward frictional force of 60N. The mass of the cyclist and bicycle is 100kg and bicycle is initially at rest. A. Calculate the speed of the bicycle after 12seconds B. what time will it take the bicycle to slow down and stop C. calculate the total distance traveled during the whole journey

OpenStudy (adi3):

@Owlcoffee @ganeshie8

OpenStudy (adi3):

@almaa @alekos

ganeshie8 (ganeshie8):

use newton's second law : \[F_{net}=ma\]

OpenStudy (adi3):

we need to find the speed

OpenStudy (welshfella):

Yes. and net force = applied force - frictional force

OpenStudy (adi3):

no it is ma

ganeshie8 (ganeshie8):

Once you have acceleration, you can find the speed using equations like \(v=u+at\) etc

OpenStudy (adi3):

the net force is 20N

OpenStudy (adi3):

u is 0

OpenStudy (adi3):

so v = 20*12

ganeshie8 (ganeshie8):

right, so 20N is acting on the cyclist-bicycle system. find the acceleration produced on the system by that net force

OpenStudy (adi3):

20/100 = 0.2

OpenStudy (adi3):

v= 0.2*12

ganeshie8 (ganeshie8):

looks good. that means, 20N force acting on the cylcist-bicycle system for 12 seconds, increases the speed from 0 m/s to 2.4 m/s

OpenStudy (adi3):

yes for b

ganeshie8 (ganeshie8):

quesiton b is not clear

OpenStudy (adi3):

ok, c

ganeshie8 (ganeshie8):

c makes no sense too

OpenStudy (adi3):

it does

ganeshie8 (ganeshie8):

is it given that the cyclist has stopped applying force on the pedals at 12th second ?

OpenStudy (adi3):

isnt b 12seconds

ganeshie8 (ganeshie8):

again, both parts b and c make no sense because the question doesn't give any info on when/how the breaks are applied

ganeshie8 (ganeshie8):

If you want to show some work to your teacher, you may assume that the force on pedals was removed at 12th second and the bicycle slows down and stops due to friction

OpenStudy (adi3):

it says if after 12seconds the cyclist stops padeling so what will happen to the motin

OpenStudy (adi3):

i did not put that bc i knew the answer

OpenStudy (adi3):

then it aks the b question

ganeshie8 (ganeshie8):

lol okay, you should put the complete question

OpenStudy (adi3):

sorry man

OpenStudy (adi3):

i think that the fiction increases

ganeshie8 (ganeshie8):

to my limited knowledge, friction disappears as soon as the cyclist removes the force on pedals. friction acts only when the bicycle accelerates... so, strictly speaking, friction cannot stop the bicycle because the friction disappears after the cyclist stops pedalling. i could be completely wrong in this matter, lets check with experts @mathmate @UnkleRhaukus

OpenStudy (adi3):

thats why it stops because of the surface fiction increasing

OpenStudy (adi3):

the fiction increases for sure

OpenStudy (adi3):

i think so

OpenStudy (adi3):

@UnkleRhaukus

ganeshie8 (ganeshie8):

If you don't apply breaks on wheels, and if there is no air drag, then i guess bicycle should never stop because the bicycle is under pure rolling(relative velocity between road and wheels is 0)

OpenStudy (adi3):

air drag is also a fiction

OpenStudy (adi3):

i think the fiction increases

OpenStudy (adi3):

right @UnkleRhaukus

ganeshie8 (ganeshie8):

are you saying 60N is the air resistance ? that makes no sense too because air resistance must depend on the velocity of bicycle, it is never constant..

OpenStudy (adi3):

no its friction while you cycle

ganeshie8 (ganeshie8):

friction from road?

OpenStudy (adi3):